alones (alones):

@jim_thompson5910 can u help me solve this one? :3

alones (alones):

Solve\[ 3x^2+7x-14=0 \]by completing the square. What are the solutions?

jimthompson5910 (jim_thompson5910):

notice how the equation is in the form ax^2 + bx + c = 0

jimthompson5910 (jim_thompson5910):

where in this case a = 3 b = 7 c = -14

jimthompson5910 (jim_thompson5910):

plug a & b into this formula h = -b/(2a) and tell me what h is

jimthompson5910 (jim_thompson5910):

let me know if you're stuck @AloneS

alones (alones):

Okay so like this? \[\frac{ -7 }{ 2(3)}\] and my answer is h= -\(\frac{ 7 }{ }\) @jim_thompson5910

alones (alones):

*-*\[\frac{ 7 }{6}\]

jimthompson5910 (jim_thompson5910):

h = -7/6, yes

jimthompson5910 (jim_thompson5910):

now plug x = -7/6 into the equation y = 3x^2+7x-14 to find the value of y

alones (alones):

Uhh i'm sorry, btu how would i do that?

jimthompson5910 (jim_thompson5910):

replace every x with -7/6, then use a calculator to compute

alones (alones):

Like this? \[-\frac{ 7 }{ 6 } \] =3x^2+7x-14

alones (alones):

Uhh I’m sorry but I don’t get it still.

jimthompson5910 (jim_thompson5910):

\[\Large y = 3x^2 + 7x - 14\] \[\Large y = 3{\color{red}{x}}^2 + 7{\color{red}{x}} - 14\] \[\Large y = 3\left({\color{red}{-\frac{7}{6}}}\right)^2 + 7\left({\color{red}{-\frac{7}{6}}}\right) - 14\] Hopefully that clears things up?

OpenStudy (obaybk3900):

yea that's right jim_thompson5910

alones (alones):

Wait than what do i do next @jim_thompson5910 :3